OpenStudy (anonymous):

Evaluate the triple integral: Integral from (0 to pi/2) integral from (0 to 1) integral from (0 to sqrt(1-r^2)) of r * sqrt(r^2+z^2) dz dr dtheta

zepdrix (zepdrix):

Just formatting it for you :) easier to read this way. \[\Large\rm \int\limits_0^{\pi/2}\int\limits_0^1\int\limits_0^{\sqrt{1-r^2}}~r\sqrt{r^2+z^2}~dz~dr~d\theta\]

OpenStudy (anonymous):

@zepdrix thank you :)

OpenStudy (anonymous):

@zepdrix help him, please!! I don't know it. :)

zepdrix (zepdrix):

Hmm our innermost integral seems a lil tough since we don't have a z outside the root :( \[\Large\rm \int\limits_0^{\sqrt{1+r^2}} r\sqrt{r^2+z^2}dz\]I guess we'll have to apply a trig substitution.\[\Large\rm z=r \tan \phi\]\[\Large\rm dz=r \sec^2 \phi ~d \phi\]Changing our integral to,\[\Large\rm \int\limits r^3 \sec^3\phi ~d\phi\]Something like that? :d Maybe there's an easier approach to this, I'm not seeing it though.

OpenStudy (anonymous):

@OOOPS thank you soooo much :) I've been stuck on this problem for a really long time